OpenStudy (anonymous):

How do you unfoil 16x^2 +48xy + 36y^2

OpenStudy (anonymous):

Through factoring.

OpenStudy (anonymous):

I don't understand factoring

OpenStudy (kinggeorge):

For factoring trinomials, it's largely guess and check. However, you'll want to notice that \(16x^2=(4x)^2\) and \(36y^2=(6y)^2\). A natural first choice for a possible factorization would then be\[(4x+6y)^2\]As it turns out, we we FOIL this again, we get \[16x^2+24xy+24xy+36y^2=16x^2+48xy+36y^2\]So we know we've found the correct answer.

OpenStudy (anonymous):

OHHH I get it! Thank you soo much

OpenStudy (kinggeorge):

You're welcome.

OpenStudy (nali):

u should close ur question so no one would mistakly open it